OpenStudy (anonymous):

Find two consecutive whole numbers that sqrt43 lies between.

OpenStudy (amistre64):

run thru your perfect squares ....

OpenStudy (anonymous):

what

OpenStudy (amistre64):

0^2 1^2 2^2 3^2 and so on; you should then be able to see where 43 falls then

OpenStudy (anonymous):

\[\begin{matrix} 6 & \sqrt{36} \\? & \sqrt{43} \\7 & \sqrt{49} \end{matrix}\]